OpenStudy (anonymous):

A bowling ball with a mass of 4.5 kg travels at a velocity of 37 m/s for 2.5 s until it is stopped at the end of the lane by the ball return. What additional information is required to determine the weight of the bowling ball? A. the speed at which the ball return captured the ball B. the amount of friction from the bowling lane acting on the bowling ball C. the acceleration due to gravity acting on the bowling ball D. the force the bowler applied to the bowling ball

OpenStudy (anonymous):

D, the force the bowler applied to the bowling ball, because F = ma, a = v/t, and m = F/a